How far should the two circular conducting plates each of radius 20/√π cm be kept in vacuum to constitute a parallel plate capacitor of capacitance 40 of?

Grade:12

1 Answers

Khimraj
5 years ago
Area of circular plate
A = \pi r^{2} = 400 cm*cm = 0.04 m*m
Capacitance C = A\varepsilon0/d
So d = 0.04\varepsilon0/40 =  8.85*10-15 m
Hope it clears.